OpenStudy (anonymous):

Hector and Yolanda buy a plasma TV from Sparky’s Electronics on an installment plan (simple interest add on loan). They will pay the TV off in 3 years, paying 19.25% interest. a. If the TV cost $1299.00, what will their monthly payments be? b. If they make an $800 down payment up front, what will their monthly payments be?

OpenStudy (anonymous):

the interest is 1299.00 * 0.1925 = 250.0575 the total price is 1299+250.0575 =1549.0575 the pay per month is 1549.0575 / (3*12) = 43.029375 you can do the same to the second one

OpenStudy (anonymous):

for part a I did.. 1299(0.1925)3=750.1725 750.1725+1299=2049.1725 2049.1725/36= 56.92 a month

OpenStudy (anonymous):

Interest=Prt

OpenStudy (anonymous):

interest per year or it for all 3 years

OpenStudy (anonymous):

doesn't say..

OpenStudy (anonymous):

Then for part B I did. 1299-800= $499 $499(0.1925)3=288.1725 288.1725+499=787.1725 787.1725/36= $21.87 a month

OpenStudy (anonymous):

why you multiply by 3 in calculating the interest ?

OpenStudy (anonymous):

I=prt Interest=(Principal)(Rate(Time) the time is in 3 years.

OpenStudy (anonymous):

yeah i think it's wright

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

u r welcome